# Laser Resonance Chromatography of 229Th3+ in He: an ab initio investigation

### v1

@@ I Introduction
The atomic structure of 229Th has sparked increasing attention over the last decades Müller et al. (2018); Fritzsche (2019). This isotope exhibits a low-lying and relatively long-lived nuclear transition from the 229mTh isomer state at about 8.338(24) eV, which may be utilized in an optical nuclear clock of unprecedented precision and stability Thirolf et al. (2024); Beeks et al. (2021); Dzuba and Flambaum (2023a); Seiferle et al. (2017); Dzuba and Flambaum (2023b); Wense et al. (2016); Müller et al. (2018); Seiferle et al. (2019); Peik and Tamm (2003). In particular, this clock could be used to investigate physics beyond the standard model, such as the space-time variation of the fundamental constants, violation of the Lorentz invariance and Einstein equivalence principle and search for dark matter fields Dzuba and Flambaum (2023a, b); Müller et al. (2018). Although 229Th can be obtained in macroscopic quantities for spectroscopy purposes, the current challenge in this field is to precisely determine the nuclear clock transition so that it can be probed with laser radiation. Once the transition is found, advanced optical clock techniques could be used to stabilize the laser frequency and control the nuclear transition. However, investigation of the clock transition is challenged by competitive nuclear relaxation channels, which in turn depend on the electronic surrounding of the nucleus Seiferle et al. (2017). Internal conversion is the dominant nuclear relaxation channel in neutral Th and leads to the emission of one electron, that is favored by the relatively small ionization potential of the atom (6.3 eV) Seiferle et al. (2017). Ionic Th has thus been considered a more promising candidate as a nuclear clock compared to the neutral, as increasing charge states raise the ionization potential and make the internal-conversion channel energetically forbidden Seiferle et al. (2017). Th3+ is nowadays considered the most promising candidate for a Th-based nuclear clock. This charge state possesses a relatively simple electronic structure that enables closed two- and three-level systems for laser-probing and cooling at the same time Block et al. (2021), and exhibits a laser-accessible 7s 2S1/2 metastable state slightly above 22000 cm-1 Safronova et al. (2006) and with 1s lifetime von der Wense et al. (2018) immune to field-induced frequency shifts von der Wense et al. (2018); Block et al. (2021). Due to these features, laser-pumping schemes were proposed to populate the 7s 2S1/2 electronic state and exploit it to monitor the nuclear clock transition Peik and Tamm (2003).
@@
In this work, we simulated the chromatographic step of an LRC experiment on 229Th3+ drifting in a helium gas, by, first, computation of accurate ion-atom interaction potentials for the two aforementioned electronic states and, second, simulation of the state-specific reduced (standard) ion mobilities over a wide range of gas temperatures and external electric field strengths, to determine the operating temperatures and field strengths that ensure state-separation. Our work is structured as follows: we detail the computational approach and methodology in Section II, while Section III is devoted to the discussion of the interaction potentials and ion mobilities computed according to Section II. Finally, the conclusions of our investigation are reported in Section IV.
@@ II.1 Ion-atom interaction potentials
The electronic structure was computed in two steps; first, Dirac-Hartree-Fock calculations were carried out on the closed-shell Th4+. The resulting wavefunction was used as the reference wavefunction for the subsequent Fock space coupled cluster calculations with iterative single and double excitations (FSCCSD) Kaldor (1991); Visscher, L. et al. (2001) calculation on the (0,1) sector. We correlated all the spinors from $4f$ up to those virtual spinors with energies below 30 $E_{h}$, with the inclusion of the $1s$ spinors of He. The model space spanned the Th $5f$, $6d$ and $7s$ spinors, for a total of thirteen spin-up and spin-down spinors. To these spinors one electron was added, whereas the remaining empty spinors below 30 $E_{h}$ were treated as virtuals. By this approach, we computed the interaction potential in a range of interatomic distances from 1.0 up to 15 Å. The interaction potentials were corrected for the basis-set superposition error by means of the procedure described by Boys and Bernardi Boys and Bernardi (1970). Eventually, we extrapolated the interaction potentials thus retrieved with the aforementioned finite basis set to the complete basis set (CBS) limit Peterson, K.A. et al. (1994), by means of the Riemann $\zeta$-formula of Lesiuk and Jeziorski Lesiuk and Jeziorski (2019). In order to assess the accuracy of our calculations in the modelling of the electron correlation, we checked the extrapolated potentials for compliance with the ion-dipole asymptotic behavior in their long-range tail,
@@ II.2 Ion mobilities
Transport of an open-shell ion undergoes polarization and inelastic collision processes. In this work, we neglect collision-induced transitions between different terms of the ion, while we account for inelastic fine-structure transitions, which involve the projections of the ion total angular momentum $J$ along the interaction axis, i.e. $\Omega$. The scattering on multiple interaction potentials thus reduces to a single-channel event, under the assumption that two different extremes approximate the collision dynamics.
@@ III.1 Ion-atom interaction potentials
We list the spectroscopic and range ($a$) parameters in Table 1.
In the short-range region before 3.5 Å, the interaction potential for the metastable $7s$ ${}^{2}S_{1/2}$ state is significantly more repulsive compared to the interaction potential for the ground $5f$ ${}^{2}F_{5/2}$ state. We ascribe this behavior to the ion’s 7s orbital, that is oriented along the interaction axis as well as the He’s 1s orbital, and thus enhances electron-electron Coulomb repulsion. For both ion states the interaction in the well region between 2.5 and 3 $\AA$ is strong, as the triply charged ion polarizes the 1s electrons of He. Nonetheless, when the ion is in its metastable state, the attractive dipole-induction term compete with the Coulomb repulsion originated from the singly occupied 7s orbital. This competition lowers by about 700 cm-1 the dissociation energy for the ion in the metastable state with respect to the analog for the ion in the ground state. For the ion ground state, the interaction is slightly anisotropic, with $\Omega$ = 5/2 and 3/2 giving rise to 10% and 7% deeper potential wells, respectively, compared to the $\Omega$ = 1/2 counterpart. We attribute this interaction anisotropy to a prolate zz-component of the permanent electric quadrupole moment associated with the Th3+ $5f$ electron. This quantity is largest for $\Omega$ = 5/2, significant for $\Omega$ = 3/2, but negligible for $\Omega$ = 1/2 (see Bellert and Breckenridge (2002) for analogous discussions).
@@
Figure 1: Graphical representations of the FSCCSD interaction potentials of the Th3+-He system calculated for the ground state $5f$ 2F5/2 and the metastable state, $7s$ 2S1/2. Counterpoise energy corrections are also included in the interaction potential. The calculated average interaction potential for the ground state is also shown (dashed line). For clarity the potentials are normalized to the same dissociation limit.
@@ III.2 Ion mobilities
Figure 2: a) Reduced zero-field mobilities of the 229Th3+-He dimer in the ground $5f$ 2F5/2 state (computed according to the isotropic and anisotropic approximations from Equation 5 and Equation 6, respectively) and metastable $7s$ 2S1/2 state; b) zero-field mobilities associated with the $\Omega$ = 1/2, 3/2, 5/2 projections of the ion’s ground state compared with the average anisotropic mobility.
@@
Figure 4: a) Reduced mobilities of the 229Th3+-He dimer in the ground $5f$ 2F5/2 state (full lines) and metastable $7s$ 2S1/2 state (dashed lines) as a function of the reduced electric field; b) relative difference between the reduced zero-field mobilities of the ground and metastable states as a function of the reduced electric field. The reduced ion mobilities are reported for four operating temperatures: 100 K (blue lines), 200 K (green lines), 300 K (orange lines) and 400 K (red lines).
@@ V Acknowledgements
We gratefully acknowledge high-performance computing (HPC) support, time and infrastructure from the Ara cluster at the Scientific Computing and Data Science Centre of the Friedrich-Schiller University in Jena, Germany. G.V. is also grateful to K.A. Peterson, A. I. Bondarev and Z.-W. Wu for suggestions and support. M.L. acknowledges funding from the European Research Council (ERC under the European Union’s Horizon 2020 Research and Innovation Programme (Grant Agreement No. 819957).

### v2

@@ I Introduction
The atomic structure of 229Th has sparked increasing attention over the last decades Müller et al. (2018); Fritzsche (2019). This isotope exhibits a low-lying and relatively long-lived nuclear transition from the 229mTh isomer state at about 8.338(24) eV, which may be utilized in an optical nuclear clock of unprecedented precision and stability Thirolf et al. (2024); Beeks et al. (2021); Dzuba and Flambaum (2023a); Seiferle et al. (2017); Dzuba and Flambaum (2023b); Wense et al. (2016); Müller et al. (2018); Seiferle et al. (2019); Peik and Tamm (2003). In particular, this clock could be used to investigate physics beyond the standard model, such as the space-time variation of the fundamental constants, violation of the Lorentz invariance and Einstein equivalence principle and search for dark matter fields Dzuba and Flambaum (2023a, b); Müller et al. (2018). Although 229Th can be obtained in macroscopic quantities for spectroscopy purposes, the current challenge in this field is to precisely determine the nuclear clock transition so that it can be probed with laser radiation. Once the transition is found, advanced optical clock techniques could be used to stabilize the laser frequency and control the nuclear transition. However, investigation of the clock transition is challenged by competitive nuclear relaxation channels, which in turn depend on the electronic surrounding of the nucleus Seiferle et al. (2017). Internal conversion is the dominant nuclear relaxation channel in neutral Th and leads to the emission of one electron, that is favored by the relatively small ionization potential of the atom (6.3 eV) Seiferle et al. (2017). Ionic Th has thus been considered a more promising candidate as a nuclear clock compared to the neutral, as increasing charge states raise the ionization potential and make the internal-conversion channel energetically forbidden Seiferle et al. (2017). Th3+ is nowadays considered the most promising candidate for a Th-based nuclear clock. This charge state possesses a relatively simple electronic structure that enables closed two- and three-level systems for laser-probing and cooling at the same time Block et al. (2021), and exhibits a laser-accessible 7s 2S1/2 metastable state slightly above 22000 cm-1 Safronova et al. (2006) and with 1s lifetime von der Wense et al. (2018) immune to field-induced frequency shifts von der Wense et al. (2018); Block et al. (2021). Due to these features, Peik and Tamm Peik and Tamm (2003) proposed a two-step laser-pumping scheme to populate the 7s 2S1/2 electronic state: in the first step, the ground state is excited to the intermediate short-living 6d 2D3/2 state, from which the metastable state is populated in the second step.
@@
In this work, we simulated the chromatographic step of an LRC experiment on 229Th3+ drifting in a helium gas, by, first, computation of accurate ion-atom interaction potentials for the two aforementioned electronic states and, second, simulation of the state-specific reduced (standard) ion mobilities over a wide range of gas temperatures and external electric field strengths, to determine the operating temperatures and field strengths that ensure state-separation. We did not consider interaction potentials and reduced ion mobilities associated with the ion in the low-lying 5f 2F7/2, and intermediate 6d 2D3/2 and 6d 2D5/2 states, since these are too short-living to survive the spectroscopic step of the experiment and, thus, do not contribute to the subsequent chromatographic step. Our work is structured as follows: we detail the computational approach and methodology in Section II, while Section III is devoted to the discussion of the interaction potentials and ion mobilities computed according to Section II. Finally, the conclusions of our investigation are reported in Section IV.
@@ II.1 Ion-atom interaction potentials
The electronic structure was computed in two steps; first, Dirac-Hartree-Fock calculations were carried out on the closed-shell Th4+. The resulting wavefunction was used as the reference wavefunction for the subsequent Fock space coupled cluster calculations with iterative single and double excitations (FSCCSD) Kaldor (1991); Visscher, L. et al. (2001) calculation on the (0,1) sector. To model the (0,0) sector, we correlated all the spinors from $4f$ up to those virtual spinors with energies below 30 $E_{h}$, with the inclusion of the $1s$ spinors of He. The model space to generate the (0,1) sector spanned the Th $5f$, $6d$ and $7s$ spinors, for a total of thirteen Kramers pairs. To these, one electron was added, whereas the remaining empty orbitals below 30 $E_{h}$ were treated as virtuals. By this approach, we computed the interaction potential in a range of interatomic distances from 1.0 up to 15 Å. The interaction potentials were corrected for the basis-set superposition error by means of the procedure described by Boys and Bernardi Boys and Bernardi (1970). Eventually, we extrapolated the interaction potentials thus retrieved with the aforementioned finite basis set to the complete basis set (CBS) limit Peterson, K.A. et al. (1994), by means of the Riemann $\zeta$-formula of Lesiuk and Jeziorski Lesiuk and Jeziorski (2019). In order to assess the accuracy of our calculations in the modelling of the electron correlation, we checked the extrapolated potentials for compliance with the ion-dipole asymptotic behavior in their long-range tail,
@@ II.2 Ion mobilities
Transport of an open-shell ion undergoes polarization and inelastic collision processes. In this work, we neglect collision-induced transitions between different terms of the ion, while we account for inelastic fine-structure transitions, which involve the projections of the ion total angular momentum $J$ along the interaction axis, i.e. $\Omega$. Indeed, this approximation holds for collisions occurring at moderate electric field strengths, such as those considered in this work, while strong-field regimes may imply non-negligible collision-induced transitions between different terms of the ion, such as collisional quenching of excited states. The scattering on multiple interaction potentials thus reduces to a single-channel event, under the assumption that two different extremes approximate the collision dynamics.
@@ III.1 Ion-atom interaction potentials
The use of the Morse potential function to fit the short-range region of heavy ion-rare gas interaction potentials is motivated by a previous work by Ramanantoanina et al. Ramanantoanina et al. (2023), where the spectroscopic parameters for the Lr+-He interaction potentials for the ground and metastable states were retrieved with reasonable accuracy. We list the spectroscopic and range ($a$) parameters in Table 1.
In the short-range region before 3.5 Å, the interaction potential for the metastable $7s$ ${}^{2}S_{1/2}$ state is significantly more repulsive compared to the interaction potential for the ground $5f$ ${}^{2}F_{5/2}$ state. We ascribe this behavior to the ion’s 7s orbital, that is oriented along the interaction axis as well as the He’s 1s orbital, and thus enhances electron-electron Coulomb repulsion. For both ion states the interaction in the well region between 2.5 and 3 $\AA$ is strong, as the triply charged ion polarizes the 1s electrons of He; in particular, the interaction potential for the $\Omega$ = 5/2 projection of the ground state features a dissociation energy $D_{e}$ of roughly 2207 cm-1, due to the strong ion-neutral induction interaction. This value is indeed in line with analogous spectroscopic parameters computed for the ground-state interaction potentials of multiply charged ions interacting with He, such as Mg2+ (7574 cm-1) and Ca2+ (1240 cm-1). However, when the ion is in its metastable state, the attractive dipole-induction term compete with the Coulomb repulsion originated from the singly occupied 7s orbital. This competition lowers by about 700 cm-1 the dissociation energy for the ion in the metastable state with respect to the analog for the ion in the ground state. For the ion ground state, the interaction is slightly anisotropic, with $\Omega$ = 5/2 and 3/2 giving rise to 10% and 7% deeper potential wells, respectively, compared to the $\Omega$ = 1/2 counterpart. We attribute this interaction anisotropy to a prolate zz-component of the permanent electric quadrupole moment associated with the Th3+ $5f$ electron. This quantity is largest for $\Omega$ = 5/2, significant for $\Omega$ = 3/2, but negligible for $\Omega$ = 1/2 (see Bellert and Breckenridge (2002) for analogous discussions).
@@
Figure 1: Graphical representations of the FSCCSD interaction potentials of the Th3+-He system calculated for the ground state $5f$ 2F5/2 and the metastable state, $7s$ 2S1/2. Counterpoise energy corrections are also included in the interaction potential. In particular, for the ground state, the interaction potentials associated with $\Omega$ = 1/2 (highest-lying full line), 3/2 and 5/2 (lowest-lying full line) are shown, along with the related average interaction potential (dashed line). For clarity the potentials are normalized to the same dissociation limit.
@@ III.2 Ion mobilities
Figure 2: a) Reduced zero-field mobilities of the 229Th3+-He dimer in the ground $5f$ 2F5/2 state (computed according to the isotropic and anisotropic approximations from Equation 5 and Equation 6, respectively) and metastable $7s$ 2S1/2 state; b) zero-field mobilities associated with the $\Omega$ = 1/2 (highest-lying full line), 3/2, 5/2 (lowest-lying full line) projections of the ion’s ground state compared with the average anisotropic mobility (dashed line).
@@
Figure 4: a) Reduced mobilities of the 229Th3+-He dimer in the ground $5f$ 2F5/2 state (full lines) and metastable $7s$ 2S1/2 state (dashed lines) as a function of the reduced electric field. The state-specific reduced ion mobilities are shown for four operating temperatures, from 100 K (lowest-lying full and dashed lines at low electric fields) to 400 K (highest-lying full and dashed lines at low electric fields), ; b) relative difference between the reduced zero-field mobilities of the ground and metastable states as a function of the reduced electric field. The reduced ion mobilities are shown for four operating temperatures, from 100 K (lowest-lying line at low electric fields) to 400 K (highest-lying line at low electric fields).
@@ IV Summary and Conclusion
Furthermore, the proposed FSCCSD-based computational approach is well suited for theoretical support of ion-mobility experiments, as ion-mobility is remarkably sensitive to dynamic electron correlation Buchachenko A.A. and Viehland (2022), which, in turn, is accurately described by methods based on the coupled cluster model. This characteristics makes the proposed computational approach scalable to the investigation of other heavy systems of interest for LRC experiments, that can be modelled by means of the FSCC level of theory. Such systems include, for instance, the ground and some low-lying excited states of Th2+ (5f6d) and Ac+ (7s2), which can be modelled by means of the (0,2)-FSCC method, and No+ (5f147s), which may be accurately modelled by means of the (0,1)-FSCC method.
@@ V Acknowledgements
We gratefully acknowledge high-performance computing (HPC) support, time and infrastructure from the Ara cluster at the Scientific Computing and Data Science Centre of the Friedrich-Schiller University in Jena, Germany. M.L. acknowledges funding from the European Research Council (ERC under the European Union’s Horizon 2020 Research and Innovation Programme (Grant Agreement No. 819957). G.V. is also grateful to K.A. Peterson, A. I. Bondarev and Z.-W. Wu for suggestions and support.